40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9952     Entry time: Wed May 14 10:04:06 2014     In reply to: 9947     Reply to this: 9973
Author: Steve 
Type: Summary 
Category: SUS 
Subject: oplev laser and temperature 

 ETMX oplev qpd gain has to be increased.

 

Attachment 1: TempRules.png  56 kB  | Hide | Hide all
TempRules.png
ELOG V3.1.3-